# Higher-Dimensional MOG dark compact object: shadow behaviour in the light of EHT observations

### v1

@@ I Introduction
Dark compact objects, from a phenomenological perspective, represent a wide range of astronomical objects, such as white dwarfs, neutron stars, and black holes. From a theoretical perspective, it is possible to predict dark compact objects within the framework of extended gravity theories and in scenarios involving particle physics beyond the standard model Cardoso:2019rvt . Gravitational waves detected by LIGO/Virgo observations recently demonstrated the occurrence of binary black hole mergers, see LIGOScientific:2017bnn and references therein. In addition, the Event Horizon Telescope (EHT) demonstrated the presence of supermassive black holes in the Milky Way and M87 galactic centers EventHorizonTelescope:2021srq ; EventHorizonTelescope:2022xqj and references therein. Thus, it is not surprising that new species of compact objects will be discovered by future developments in the fields of very long baseline interferometry (VLBI) APhaseSensitiveInterferometer:1985 and gravitational wave astronomy. However, the study of dark compact objects from a mathematical perspective is interesting, because it can assist in comprehending how and at what limits a dark compact object tends to be a black hole by increasing its compactness. As mentioned, the Event Horizon Telescope (EHT) collaboration’s pioneering achievement of constructing the first pictures of the shadows of the supermassive black holes M87* (see EventHorizonTelescope:2021srq and references therein) and Sgr A* (see EventHorizonTelescope:2022xqj and references therein) has opened a new chapter in the physics of black holes and dark compact objects and enhanced our understanding of these mysterious celestial objects. Since a black hole does not emit light, its event horizon, that is, the surface at which matter or radiation cannot pass through, is not directly observable. Instead, what we can see is the ”shadow” of a black hole, a dark region surrounded by light that results in the gravitational lensing phenomenon around the event horizon  Perlick:2004tq ; Cosmpactness:2022 ; Relativisticimage:2009 ; ImagesDistortionHypothesis:2024 ; Cosmologicalconstant:2022 ; Ishihara:2016 ; Tsukamoto:2017 ; Ishihara:2017 ; Virbhadra:2000 ; Claudel:2001 ; Virbhadra:2002 ; Virbhadra:1998 ; Tsukamotoa:2023 ; Kumaran:2023 . Many efforts have been made to improve measurements in order to obtain high-resolution images since the release of the shadow images of M87* and Sgr A*  Goddi:2016qax . To achieve the needed resolution, theoretical efforts to study black hole physics, in particular, shadow behavior in different gravitational theories, has become of increasing significance. New theoretical shadow templates for upcoming observations are provided in this regard by analytical and numerical studies and analyses of the apparent geometrical shape of different black hole spacetimes. In addition to spacetime characteristics  Johannsen:2010ru ; Cunha:2017eoe and the observer’s position, the black hole parameters; mass, electric charge, and angular momentum determine the shadow’s shape and size deVries:1999tiy . The shadow shape for non-rotating black holes is a perfect circle. Rotating black holes can have non-trivial shadow patterns due to the angular momentum parameter  Chandrasekhar:1984siy .
@@
Even though Albert Einstein’s General Theory of Relativity (GR) has made significant progress in explaining observations and predicting astounding events, it is still unable to fully explain gravitational interaction and the events that correspond with it in the universe. Some instances of the shortcomings of GR are the reproduction of the rotation curves of nearby galaxies Sofue:2000jx ; Sofue2016 , mass profiles of galaxy clusters Ettori2013 ; Voigt2006 , inherent singularities at the center of black holes, etc. Furthermore, the positively accelerated expansion of the Universe at late time is explained by GR only if the cosmological constant factor $\Lambda$ is present Weinberg1989 ; Peebles2003 . Restructuring the geometric part of General Relativity (GR) using various ways is an intriguing reform strategy. One such approach is the so-called MOdified Gravity (MOG), a Scalar-Tensor-Vector (STVG) theory developed and proposed by John W. Moffat to describe gravitational interaction Moffat2006 . The gravitational effects of spacetime in the MOG setup are expressed by a massive vector field $\phi_{\mu}$ in addition to three scalar fields: the mass of the vector field $\tilde{\mu}$, the effective gravitational constant $G$, and the vector field coupling $\xi$. In addition to being compatible with cosmological observational data Davari2021 and Planck 2018 data  Moffat2021a , the MOG theory has achieved various successes in explaining astrophysical phenomena. These include explaining the dynamics of galaxy clusters without dark matter and the rotation curves of many galaxies  Brownstein2006a ; Brownstein2006b ; Brownstein2007 ; Moffat2013 ; Moffat2014 ; Moffat2015a . Furthermore, the framework of MOG theory is currently experiencing the release of a number of black hole solutions, including non-rotating and rotating ones  Moffat2015b , even with extra dimensions Cai2021 , cosmological solutions Roshan2015 ; Jamali2018 ; Davari2021 , and non-stationary solutions for inhomogeneity distributions of mass–energy in spacetime  Perez2019 . Additionally, a great deal of theoretical and observational work has been done to comprehend the characteristics of the MOG theory and how it functions in various contexts  Moffat2021b ; Guo2018 ; Moffat2020 ; Mureika2016 ; Saghafi2021 ; Saghafi:2021wzx ; Perez:2017spz ; Shukla:2022sti ; Hussain:2015cga ; Sharif:2017owq ; Lee:2017fbq ; Kolos:2020ykz ; Hu:2022lek ; John:2019was . Ref.  Moffat:2018jmi offers an interesting exploration of the solution characterizing the regular rotating and non-rotating MOG dark compact object. Ref.  Sau:2022afl examines the shadow behavior of the regular rotating and non-rotating MOG dark compact object.
@@
The notion of extra dimensions was initially put forward by Nordstrom in 1914  Nordstrom:1914ejq . His theory states that by treating four-dimensional spacetime as a surface in a five-dimensional spacetime, one may unify the gravitational and electromagnetic fields. Currently, the primary motivations of the vast amount of research on the extra dimensions are the quantization of gravitational interaction, the unification of gauge and gravitational interactions of elementary particles, the Higgs mass hierarchy problem, and the cosmological constant problem. This is why the Kaluza-Klein (KK) theory Kaluza:1921tu ; Klein:1926tv unifies electromagnetic or even non-Abelian gauge fields that characterize weak and strong interactions with gravitational interaction by introducing a compact space constructed by compact extra dimensions with a specific compactification scale. Additionally, as a well-known candidate theory for quantum gravity, the string theory (M-theory) has ten or eleven compact dimensions Witten:1995ex ; Schwarz:1995jq . There are ideas for large additional dimensions of the order of millimeter in addition to these compact extra dimensions with the extension up to the order of the Planck length. The Arkani-Hamed-Dimopoulos-Dvali (ADD) braneworld model Arkani-Hamed:1998jmv ; Arkani-Hamed:1998sfv has provided a novel path for the study of extra dimensions by using large extra dimensions to solve the Higgs mass hierarchy problem. It is important to remember that the striking aspect of these large extra dimensions is that future accelerator, astrophysical, and even tabletop studies will be able to measure their effects. It is surprising to learn that string theory can incorporate the ADD model  Antoniadis:1998ig . The Randall-Sundrum (RS) braneworld model Randall:1999ee ; Randall:1999vf proposes the use of warped extra dimensions in addition to compact and large extra dimensions to solve the Higgs mass hierarchy problem. Apart from the aforementioned extra dimensions, there exist theories with unlimited volume extra dimensions such as the Dvali-Gabadadze-Porrati (DGP) braneworld scenario  Dvali:2000hr , wherein the spacetime is not four-dimensional even at very low energies and the extra dimensions are neither wrapped nor compact. These theories, which alter gravity at large distances because the existence of infinite volume additional dimensions, are possibilities for solving the cosmological constant problem Dvali:2000xg ; Dvali:2002pe . Refs. Gabadadze:2003ii ; Shifman:2009df ; Perez-Lorenzana:2005fzz ; Rubakov:2001kp provide some thorough evaluations of higher-dimensional models. A variety of techniques and strategies have been used within the context of black hole physics to expand and study alternative black hole models in arbitrary dimensions Emparan:2008eg ; Kanti:2004nr . One such technique is the Tangherlini method Tangherlini:1963bw , which generalizes the Schwarzschild solution to $n$ dimensions.
@@
In high-energy or particle experiments, physicists concentrate on detecting extra dimensions. Prospective instruments for investigating the impact of strong gravity regimes corresponding to higher-dimensional black holes and extra dimensions are the Large Hadron Collider (LHC) at CERN and future colliders Allanach:2002gn ; Agashe:2006hk ; Franceschini:2011wr ; Deutschmann:2017bth ; Strominger:1996sh ; Harris:2004xt ; Antoniadis:2000vd . Furthermore, evidence for the existence of extra dimensions comes from spectroscopic tests Zhou:2014xbw ; Luo:2006ck ; Luo:2006ad , the presence of hydrogen atoms in higher dimensions Burgbacher:1999sha ; Caruso:2012daf ; Shaqqor:2009cha , and theories to solve the proton radius puzzle Wang:2013fma ; Dahia:2015bza ; Zhi-gang:2007swh . However, the detection of gravitational waves (GW) by the LIGO/Virgo collaborations LIGOScientific:2016aoc and the aforementioned images obtained by the EHT collaborations are two recent successes in the search for black hole strong field regimes. Some traces of additional dimensions exist in the GW detection, with specific details regarding the associated amplitude and fluctuation mode dynamics. As a result, a lot of research has been done to uncover this physics Cardoso:2016rao ; Yu:2016tar ; Visinelli:2017bny ; Kwon:2019gsa ; for a thorough analysis, refer to Ref. Yu:2019jlb . EHT has now opened up novel opportunities for the search for extra dimensions. The authors of recent seminal publications Vagnozzi:2019apd ; Belhaj:2020mlv ; Tang:2022hsu discovered significant limitations on warped and compact extra dimensions within the RS model and M-theory, derived from EHT observations. Thus, as we aim to do in this study, one can use the EHT data to investigate all kinds of extra dimensions in general and see if they can be detected. In this sense, it appears that extra dimensions have an impact on black hole shadows by decreasing the size of the shadow in a variety of black hole models and gravitational theories Amarilla:2011fx ; Eiroa:2017uuq ; Papnoi:2014aaa ; Singh:2017vfr ; Amir:2017slq ; Belhaj:2020rdb . Nevertheless, there is still much to learn about the precise impact of extra dimensions on black hole shadows, which is the subject of ongoing research. It appears that there is a greater chance of future detection of the effects of large and infinite volume extra dimensions.
@@
Furthermore, due to extra degrees of freedom that arose from extended theories of gravity, black hole shadow size and shape may vary. Thus, determining the size and shape of black hole shadows could help evaluate black hole metrics parameters and test alternative theories of gravity. The question of black hole shadow has been the subject of a great deal of research to determine whether and how degrees of freedom—derived from extended theories of gravity other than black hole parameters—affect the behavior of the shadow Perlick:2021aok . Here are a few instances: Black hole shadows in Einstein-Maxwell-dilaton gravity Amarilla:2013sj ; Wei:2013kza , in Chern-Simons modified gravity Amarilla:2010zq , the shadow behavior of the Kerr-Newman family of solutions of the Einstein-Maxwell equations is studied in Refs. Bardeen:1973tla ; Takahashi:2004xh ; the shadow of a black hole with NUT-charges Chakraborty:2013kza ; Grenzebach:2014fha ; the apparent shape of the Sen black hole  Hioki:2008zw ; Dastan:2016bfy ; Younsi:2016azx ; shadow behavior of the non-commutative geometry inspired and quantum-corrected black holes  Wei:2015dua ; Sharif:2016znpx ; Saghafi:2022pmey ; shadow behavior of Einstein-Born-Infeld black holes  Atamurotov:2015xfa ; shadow behavior of Ayon-Beato-Garcia black hole and also, rotating Hayward and rotating Bardeen regular black holes  Abdujabbarov:2016hnw and hairy black holes Cunha:2015yba ; Cunha:2016bjh ; shadows of colliding and multi-black holes Nitta:2011nin ; Yumoto:2012kz ; shadow behavior of rotating black holes in $f(R)$ gravity Dastan:2016vhb , conformal Weyl gravity  Mureika:2016efo , and Einstein-dilaton-Gauss-Bonnet black holes  Cunha:2016wzk ; Shadow behavior of wormholes and naked singularities Nedkova:2013msa ; Ohgami:2015nra ; Ortiz:2015rma , as well as chaotic shadow of a non-Kerr rotating compact object with quadrupole mass moment and a magnetic dipole  Wang:2018eui ; Wang:2017qhh , and black holes with exotic matter  Tinchev:2015apf ; Abdujabbarov:2015pqp ; Singh:2017xle ; Huang:2016qnl .
@@
This paper is organized as follows. In section II we first introduce the Modified Gravity theory and then introduce the line element of the higher-dimensional dark compact object in this theory. In section III we study the shadow behavior, energy emission rate, and deflection angle of the STVG dark compact object in MOG theory with extra dimensions, and we analyze the impact of parameter $\alpha$ and dimensions on the shadow and deflection angle of the dark compact object. In section IV we constrain the parameter $\alpha$ by comparing the shadow size of the suppermassive black hole M87* with the shadow size of the STVG dark compact object. In section V we examin the issue of acceleration bounds in higher-dimensional MOG dark compact object’s spacetime background. Finally, Section VI is devoted to summary, conclusion and discussing about our main results.

### v3

@@ I Introduction
Dark compact objects, from a phenomenological perspective, represent a wide range of astronomical objects, such as white dwarfs, neutron stars, and black holes. From a theoretical perspective, it is possible to predict dark compact objects within the framework of extended gravity theories and in scenarios involving particle physics beyond the standard model Cardoso:2019rvt . Gravitational waves detected by LIGO/Virgo observations recently demonstrated the occurrence of binary black hole mergers, see LIGOScientific:2017bnn and references therein. In addition, the Event Horizon Telescope (EHT) demonstrated the presence of supermassive black holes in the Milky Way and M87 galactic centers EventHorizonTelescope:2021srq ; EventHorizonTelescope:2022xqj and references therein. Thus, it is not surprising that new species of compact objects will be discovered by future developments in the fields of very long baseline interferometry (VLBI) APhaseSensitiveInterferometer:1985 and gravitational wave astronomy. However, the study of dark compact objects from a mathematical perspective is interesting, because it can assist in comprehending how and at what limits a dark compact object tends to be a black hole by increasing its compactness. As mentioned, the Event Horizon Telescope (EHT) collaboration’s pioneering achievement of constructing the first pictures of the shadows of the supermassive black holes M87* (see EventHorizonTelescope:2021srq and references therein) and Sgr A* (see EventHorizonTelescope:2022xqj and references therein) has opened a new chapter in the physics of black holes and dark compact objects and enhanced our understanding of these mysterious celestial objects. Since a black hole does not emit light, its event horizon, that is, the surface at which matter or radiation cannot pass through, is not directly observable. Instead, what we can see is the ”shadow” of a black hole, a dark region surrounded by light that results in the gravitational lensing phenomenon around the event horizon  Perlick:2004tq ; Cosmpactness:2022 ; Relativisticimage:2009 ; ImagesDistortionHypothesis:2024 ; Cosmologicalconstant:2022 ; Ishihara:2016 ; Tsukamoto:2017 ; Ishihara:2017 ; Virbhadra:2000 ; Claudel:2001 ; Virbhadra:2002 ; Virbhadra:1998 ; Tsukamotoa:2023 ; Kumaran:2023 . Much attempts have been made to measure properly in order to obtain high-resolution images since the release of the shadow images of M87* and Sgr A*  Goddi:2016qax . To achieve the needed resolution, theoretical efforts to study physics of black hole, in particular, shadow behavior in different theories of gravitation, has become of increasing significance. Novel models of theoretical shadow for upcoming observations are provided in this case by numerical and analytical studies and analyses of the geometric shape of different black hole spacetimes. More than spacetime characteristics  Johannsen:2010ru ; Cunha:2017eoe and the observer’s place, the parameters of black hole; angular momentum, mass , and electric charge determine the shadow’s shape and size deVries:1999tiy . The shadow shape is a circle and a perfect one for black holes which are non-rotating. But black holes which are rotating may have got non-trivial shadow patterns due to the angular momentum parameter  Chandrasekhar:1984siy .
@@
Even though Albert Einstein’s General Theory of Relativity (GR) has made significant progress in explaining observations and predicting astounding events, it is still unable to fully explain gravitational interaction and the events that correspond with it in the universe. Some instances of the shortcomings of GR are the reproduction of the rotation curves of nearby galaxies Sofue:2000jx ; Sofue2016 , mass profiles of galaxy clusters Ettori2013 ; Voigt2006 , inherent singularities at the center of black holes, etc. Furthermore, the positively accelerated expansion of the Universe at late time is explained by GR only if the cosmological constant factor $\Lambda$ is present Weinberg1989 ; Peebles2003 . Restructuring the geometric part of General Relativity (GR) using various ways is an intriguing reform strategy. One such approach is the so-called MOdified Gravity (MOG), a Scalar-Tensor-Vector (STVG) theory developed and proposed by John W. Moffat to describe gravitational interaction Moffat2006 . The gravitational effects of spacetime in the MOG setup are expressed by a massive vector field $\phi_{\mu}$ in addition to three scalar fields: the mass of the vector field $\tilde{\mu}$, the effective gravitational constant $G$, and the vector field coupling $\xi$. In addition to being compatible with cosmological observational data Davari2021 and Planck 2018 data  Moffat2021a , the MOG theory has achieved various successes in explaining astrophysical phenomena. These include explaining the dynamics of galaxy clusters without dark matter and the rotation curves of many galaxies  Brownstein2006a ; Brownstein2006b ; Brownstein2007 ; Moffat2013 ; Moffat2014 ; Moffat2015a . Furthermore, the framework of MOG theory is currently experiencing the release of a number of black hole solutions, including non-rotating and rotating ones  Moffat2015b , even with extra dimensions Cai2021 , cosmological solutions Roshan2015 ; Jamali2018 ; Davari2021 , and non-stationary solutions for inhomogeneity distributions of mass–energy in spacetime  Perez2019 . Additionally, a great deal of theoretical and observational work has been done to comprehend the characteristics of the MOG theory and how it functions in various contexts  Moffat2021b ; Guo2018 ; Moffat2020 ; Mureika2016 ; Saghafi2021 ; Saghafi:2021wzx ; Perez:2017spz ; Shukla:2022sti ; Hussain:2015cga ; Sharif:2017owq ; Lee:2017fbq ; Kolos:2020ykz ; Hu:2022lek ; John:2019was ; Camilloni:2024 ; rosa:2023 . Ref.  Moffat:2018jmi offers an interesting exploration of the solution characterizing the regular rotating and non-rotating MOG dark compact object. Ref.  Sau:2022afl examines the shadow behavior of the regular rotating and non-rotating MOG dark compact object.
@@
The notion of higher dimensions was initially put forward by Nordstrom in 1914  Nordstrom:1914ejq . His theory states that Using four-dimensional spacetime as a surface in a five-dimensional spacetime, one may unify the gravitational and electromagnetic fields. Currently, the primary motivations of the vast amount of research on the extra dimensions are the quantization of gravitational interaction, the unification of gauge and gravitational force among elementary particles, the the cosmological constant problem, and Higgs mass hierarchy problem. This is why the Kaluza-Klein (KK) theory Kaluza:1921tu ; Klein:1926tv unifies non-Abelian gauge fields or electromagnetic that feature weak and strong forces with gravitational interaction by introducing a compact space built from compact extra dimensions with a specific compactification scale. Additionally, as a famous theory,a candidate for quantum gravity, the string theory (M-theory) has ten or eleven compact dimensions Witten:1995ex ; Schwarz:1995jq . There are some thoughts for large additional dimensions of the order of millimeter in addition to these compact extra dimensions with the extension up to the order of the Planck length. The Arkani-Hamed-Dimopoulos-Dvali (ADD) braneworld model Arkani-Hamed:1998jmv ; Arkani-Hamed:1998sfv has provided a novel path for the study of extra dimensions by using large extra dimensions to solve the Higgs mass hierarchy problem. It is important to remember that the striking aspect of these large extra dimensions is that future accelerator, astrophysical, and even tabletop studies will be able to measure their effects. It is surprising to learn that string theory can incorporate the ADD model  Antoniadis:1998ig . The Randall-Sundrum (RS) braneworld model Randall:1999ee ; Randall:1999vf proposes the use of warped extra dimensions in addition to compact and large higher dimensions to solve the Higgs mass hierarchy problem. Apart from the aforementioned extra dimensions, there exist theories with unlimited volume extra dimensions such as the Dvali-Gabadadze-Porrati (DGP) braneworld scenario  Dvali:2000hr , wherein the spacetime is not four-dimensional even at very low energies and the higher dimensions are not wrapped and not compact. These theories, which alter gravity at large distances because the existence of infinite volume additional dimensions, are possibilities for solving the cosmological constant problem Dvali:2000xg ; Dvali:2002pe . Refs. Gabadadze:2003ii ; Shifman:2009df ; Perez-Lorenzana:2005fzz ; Rubakov:2001kp provide some thorough evaluations of higher-dimensional models. A variety of techniques and strategies have been used in the context of physics of black holes to expand and study alternative solutions for black holes in arbitrary dimensions Emparan:2008eg ; Kanti:2004nr . One such technique is the Tangherlini method Tangherlini:1963bw , which generalizes the solution for Schwarzschild to $n$ dimensions.
@@
Physicists concentrate on detecting extra dimensions in high energy regimes. Prospective instruments for investigating the impact of regimes of strong gravity related to black holes with extra dimensions and extra dimensions are the Large Hadron Collider at CERN and future colliders Allanach:2002gn ; Agashe:2006hk ; Franceschini:2011wr ; Deutschmann:2017bth ; Strominger:1996sh ; Harris:2004xt ; Antoniadis:2000vd . Furthermore, evidence for the existence of extra dimensions comes from spectroscopic tests Zhou:2014xbw ; Luo:2006ck ; Luo:2006ad , the existence of hydrogen atoms in extra dimensions Burgbacher:1999sha ; Caruso:2012daf ; Shaqqor:2009cha , and theories to solve the proton radius puzzle Wang:2013fma ; Dahia:2015bza ; Zhi-gang:2007swh . However, the detection of gravitational waves (GW) by the LIGO/Virgo collaborations LIGOScientific:2016aoc and the aforementioned images obtained by the EHT collaborations are two recent successes in the search for black hole strong field regimes. Some signs of additional dimensions are in the GW detection, with specific details regarding the associated amplitude and fluctuation mode dynamics. As a result, a lot of research has been done to uncover this physics Cardoso:2016rao ; Yu:2016tar ; Visinelli:2017bny ; Kwon:2019gsa ; for a thorough analysis, refer to Ref. Yu:2019jlb . EHT has now opened up novel opportunities for the search for extra dimensions. The authors of recent seminal publications Vagnozzi:2019apd ; Belhaj:2020mlv ; Tang:2022hsu discovered significant limitations on warped and compact extra dimensions beyond the M-theory and RS model, derived from EHT observations. Thus, as we aim to do in this study, the EHT data can be used to investigate whole kinds of higher dimensions in general and notice if they can be observed. In this sense, it appears that extra dimensions have an impact on black hole shadows by decreasing the size of the shadow in a variety of black hole models and gravitational theories Amarilla:2011fx ; Eiroa:2017uuq ; Papnoi:2014aaa ; Singh:2017vfr ; Amir:2017slq ; Belhaj:2020rdb . Nevertheless, there is still much to learn about the precise impact of higher dimensions on black hole shadows, which is the subject of ongoing research.
@@
Furthermore, due to additional degrees of freedom that comes from modified theories of gravity, black hole shadow size and shape may vary. Thus, determining the size and shape of black hole shadows could help evaluate black hole metrics parameters and test alternative theories of gravity. The question of black hole shadow has been the subject of a great deal of research to determine whether degrees of freedom affect the behavior of the shadow Perlick:2021aok . Here are a few instances: Einstein-Maxwell-dilaton gravity black holes and their shadows Amarilla:2013sj ; vag:2023 ; vag1:2023 ; vag2:2023 ; vag3:2023 ; vag4:2023 ; vag5:2023 ; Wei:2013kza , in Chern-Simons modified gravity Amarilla:2010zq , the Kerr-Newman family of solutions of the Einstein-Maxwell equations and its shadow behavior is studied in Refs. Bardeen:1973tla ; Takahashi:2004xh ; the black hole with NUT-charges and its shadow Chakraborty:2013kza ; Grenzebach:2014fha ; the apparent shape of the Sen black hole  Hioki:2008zw ; Dastan:2016bfy ; Younsi:2016azx ; the non-commutative geometry inspired and quantum-corrected black holes and their shadows  Wei:2015dua ; Sharif:2016znpx ; Saghafi:2022pmey ; Einstein-Born-Infeld black holes and its shadow  Atamurotov:2015xfa ; Ayon-Beato-Garcia black hole and also, rotating Hayward and rotating Bardeen regular black holes and their shadows  Abdujabbarov:2016hnw and hairy black holes Cunha:2015yba ; Cunha:2016bjh ; colliding and multi-black holes and shadows Nitta:2011nin ; Yumoto:2012kz ; investigating shadows of rotating black holes in $f(R)$ gravity Dastan:2016vhb , conformal Weyl gravity  Mureika:2016efo , and Einstein-dilaton-Gauss-Bonnet black holes  Cunha:2016wzk ; studying Shadow of wormholes and naked singularities Nedkova:2013msa ; Ohgami:2015nra ; Ortiz:2015rma , as well as chaotic shadow of a non-Kerr rotating compact object with quadrupole mass moment and a magnetic dipole  Wang:2018eui ; Wang:2017qhh , and black holes with exotic matter  Tinchev:2015apf ; Abdujabbarov:2015pqp ; Singh:2017xle ; Huang:2016qnl .
@@
This paper is based as follows. In section II first the Modified Gravity theory is introduced and then we introduce the line element of the higher-dimensional dark compact object in this theory. In section III we study the shadow behavior, energy emission rate, and deflection angle of the STVG dark compact object in MOG theory with higher dimensions, and the impact of parameter $\alpha$ and dimensions on the shadow and deflection angle of the dark compact object is analyzed. In section IV we constrain the parameter $\alpha$ by comparing the shadow size of the suppermassive black hole M87* with the shadow size of the STVG dark compact object. In section V we examin the issue of acceleration bounds in higher-dimensional MOG dark compact object’s spacetime background. Finally, Section VI would discuss summary, conclusion and our main results.